Ship reports banner with links Home IODP-USIO daily/weekly reports

< Previous expedition | Next expedition >

IODP Expedition 328

Cascadia A-CORK Observatory

Daily Science Report for 18 September 2010

LOCATION: In transit/Esquimalt Graving Dock, Victoria, B.C.

SCIENCE UPDATE: Completed transit from Site U1364 to Victoria. First line was ashore at 0800 hr, ending Expedition 328.


Daily Science Report for 17 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The inflate bridge plug was run into the hole to 319 mbsf and set and released 4 m above the casing shoe. Five barrels of 16 ppg cement were pumped above packer forming a plug ~15 m thick. After circulating the hole with seawater, the drilling string was tripped out of the ACORK and up to the rig floor. The ship was secured for transit and underway to Victoria by 1400 hr.


Daily Science Report for 16 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The ACORK assembly continued to run in the hole to 326.7 mbsf and was landed in the re-entry cone at 00:25 hr without incident. The ROV platform was assembled and deployed, landing on the re-entry cone and released by 06:00 hr. A cone was assembled and free fall deployed over the ACORK head to facilitate re-entry. When pulling out of the ACORK the end of the drill string dislodged the funnel, which fell to the seafloor. The drill string was tripped to the surface, and the inflatable bridge plug was assembled and the drill string was tripped to the seafloor to re-entry the ACORK without the aid of the funnel. At 23:20 hr, Hole U1364A was re-entered through the ACORK and the drilling string was run into the hole reaching 99.1 mbsf by midnight.


Daily Science Report for 15 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: Running 10-3/4 in casing, attaching the umbilical, and making tubing connections continued from 150-323 m. The ACORK was picked up and attached to the top joint of casing. The BHA with mud motor, under-reamer, and running tool was made up and landed into the ACORK. Final plumbing connections were made on the ACORK, which was submerged to purge lines, purge valves were closed, and the assembly was run to 418.7 mbrf. The VIT was deployed to confirm that the bit and under-reamer extended beneath casing and then the ACORK assembly was run to the re-entry cone, which was re-entered at 19:30 hr. By midnight, the bit had reached 287 mbsf.


Daily Science Report for 14 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: Drilling Hole U1364A continued reaching the total depth of 336 mbsf at 0645 hr. After a wiper trip and displacing the hole with heavy mud, the drill string was tripped back to the rig floor. Rig up and deployment of 10-3/4 inch casing, screens, and umbilical followed with 150 m of casing, including 3 screens joints deployed by midnight.


Daily Science Report for 13 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The mud motor and under-reamer were successfully tested and the ACORK was picked up and racked in the derrick. A 14-3/4 inch drilling assembly was made up and deployed. Hole U1364A was re-enterted at 0905 hr to drill to total depth (336 mbsf). Drilling continued throughout the day and reached 260.8 mbsf by midnight.


Daily Science Report for 12 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The re-entry cone and casing was launched, tripped to seafloor, and jetted-in to 53 mbsf. The running tool was released and the drilling assembly was tripped back to rig floor. The mud motor and under-reamer were made up for deck testing.


Daily Science Report for 11 September 2010

LOCATION: Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The jet-in test was completed to 55 mbsf and the drill string was tripped back to the rig floor. Three joints of 16 inch casing and a 16 inch casing hangar were made up and landed in the re-entry cone. Total string length is 52.9 m. Made up casing jet-in assembly with a 14-3/4 bit. Completed testing of the ACORK pressure bay.


Daily Science Report for 10 September 2010

LOCATION: In transit/Site U1364 (48° 41.9964'N, 126° 52.3302'W)

SCIENCE UPDATE: The ship arrived on location at 0600 hr. After positioning on site, the drilling string was tripped to the seafloor. The Vibration Isolated Television (VIT) was deployed to visually verify tag of seafloor at 1329 meters below rig floor (mbrf), then recovered. The jet-in test was initiated and had reached 35 m below seafloor (mbsf) by midnight.


Daily Science Report for 9 September 2010

LOCATION: Ogden Point, Pier A, Victoria, B.C./Underway to Site U1364

SCIENCE UPDATE: Completed loading and unloading activities. The last line was released and the ship departed at 1606 hr and began the transit to Site U1364.


Daily Science Report for 8 September 2010

LOCATION: Ogden Point, Pier A, Victoria, B.C.

SCIENCE UPDATE: Logistical activities focused on intensive loading and offloading of equipment and supplies. Pressure transducers were mounted to the ACORK well head. Introductory meetings were conducted with Expedition 328 participants and School of Rock initiated formal activities. Estimated time of departure is ~1600 hr 9 September.


Daily Science Report for 7 September 2010

LOCATION: Ogden Point, Pier A, Victoria, B.C.

SCIENCE UPDATE: Port call activities continued. A busy public relations day included a press conference, tours for IODP-Canada, US National Research Council Committee on Scientific Ocean Drilling, and the public. School of Rock participants moved onboard.


Daily Science Report for 6 September 2010

LOCATION: Ogden Point, Pier A, Victoria, B.C.

SCIENCE UPDATE: Port call activities continued. Four public relation tours were conducted and School of Rock instructors moved onboard.


Daily Science Report for 5 September 2010

LOCATION: Ogden Point, Pier A, Victoria, B.C.

SCIENCE UPDATE: The first line was ashore at 0836 hr, ending Expedition 327 and beginning Expedition 328. After clearing customs and immigration, Expedition 327 scientists disembarked. USIO crew change and crossover was completed. Freight loading and unloading activities were initiated.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Tuesday, 08-Nov-2011 08:54:54 CST.